OpenStudy (anonymous):

Write the equation of a vertical line that goes through the point (–3, 18). can you help me do this

OpenStudy (raden_zaikaria):

y=mx + c

OpenStudy (anonymous):

i know that part but am i doing y2-y1/x2-x1

OpenStudy (raden_zaikaria):

( x,y ) is ( -3, 18 )

OpenStudy (anonymous):

im very confused can you show me.... y=mx+b is point slope formula?

OpenStudy (raden_zaikaria):

must find m with tht formula first

OpenStudy (raden_zaikaria):

wait

OpenStudy (anonymous):

how do i find the slope then?

OpenStudy (raden_zaikaria):

( -3,18) and ( -3,0 ) because that vertical line.... dont have the slope

OpenStudy (anonymous):

so what would that equation be?

OpenStudy (raden_zaikaria):

y2-y1/x2-x1 for find M

OpenStudy (anonymous):

there is no slope

OpenStudy (raden_zaikaria):

can u show me the full question ....

OpenStudy (raden_zaikaria):

no slope... m=0

OpenStudy (anonymous):

Write the equation of a vertical line that goes through the point (–3, 18). Show your work.

OpenStudy (raden_zaikaria):

yes... because that line is vertical... no slope... :)

OpenStudy (anonymous):

so am i doing it right so far?

OpenStudy (anonymous):

what you my actual equation be? it has it y=mx+b right